OpenStudy (anonymous):

Does 6^2=(5+3)*8-2 Equal 2?

OpenStudy (anonymous):

no you can tel instantly by LHS it doesn't equal 2

OpenStudy (across):

What you have here is a contradiction.

OpenStudy (anonymous):

36=8*8-2 36=64-2 Then What?

OpenStudy (across):

You are basically claiming that \(36\) is equal to \(62\), or that pears are the same thing as donkeys.

OpenStudy (anonymous):

\[\frac{ 36 }{ 36} =\frac{ 62}{ 36 } Isnt That Right?\]

OpenStudy (across):

You're basically claiming that \(1\) is equal to \(1.7\bar2\), or that pears are the same thing as donkeys.

OpenStudy (anonymous):

If You Divide 36 By 36 Then They Cross Out!

OpenStudy (across):

\(36\) divided by \(36\) is \(1\).

OpenStudy (anonymous):

No If You It Crosses Out! Nvm I'll Ask Someone Else @timo86m Can You Help Me?

OpenStudy (across):

You do realize that you're saying that \(1=1.7\bar2\), do you? :P

OpenStudy (anonymous):

No! Nvm You Dont Understand What Im Trying To Do!

OpenStudy (across):

What is it that you are trying to do?

OpenStudy (anonymous):

Im Doing It Based On Order Of Operations

OpenStudy (across):

\[6^2=8(5+3)-2\\ 36=8\cdot8-2\\ 36=64-2\\ 36=62\leftarrow\text{This is false.}\]

OpenStudy (anonymous):

Ok So Thats The Answer Or Not?

OpenStudy (across):

I personally would say that the answer is: "This is a false statement." Are you sure you copied the problem correctly?

OpenStudy (anonymous):

Yes! Im Sure!

OpenStudy (anonymous):

Im Just Going To Put A Question Mark On It Thats Why He Can Help Me With It Tommorrow!